# A Dirichlet stochastic block model for composition-weighted networks

### v1

@@ 3.2 Identifiability of parameter ordering
The DirSBM exhibits a minor non-idenfitiability issue related to the ordering of clusters and the ordering of parameters. In particular, once we obtain the final node partition and the parameter estimates using the algorithm described in Section 3.1, the ordering of rows of the estimated Dirichlet concentration matrix $\hat{\mathbf{A}}$ and of the entries of the mixing proportions vector $\hat{\boldsymbol{\theta}}$ is not guaranteed to match the ordering of clusters in the partition.
To aid understanding, consider an example with
$\mathbf{c}=(1,1,2,1,2),\hskip 2.84526pt\mathbf{A}=\begin{pmatrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{pmatrix}\text{ and }\boldsymbol{\theta}=(\theta_{1},\theta_{2}).$
For observation 1, $\mathbf{x}_{1}^{*}\sim\theta_{1}Dir(\alpha_{11},\alpha_{12},\alpha_{11},\alpha_{12})+\theta_{2}Dir(\alpha_{21},\alpha_{22},\alpha_{21},\alpha_{22})$. Now, consider the same partition of nodes with the same cluster order and reorder the rows of the parameter matrix $\mathbf{A}$ and the entries of $\boldsymbol{\theta}$, i.e.
$\mathbf{c}=(1,1,2,1,2),\hskip 2.84526pt\mathbf{A}^{\prime}=\begin{pmatrix}\alpha_{21}&\alpha_{22}\\ \alpha_{11}&\alpha_{12}\end{pmatrix}=\begin{pmatrix}\alpha_{11}^{\prime}&\alpha_{12}^{\prime}\\ \alpha_{21}^{\prime}&\alpha_{22}^{\prime}\end{pmatrix}\text{ and }\boldsymbol{\theta}^{\prime}=(\theta_{2},\theta_{1})=(\theta_{1}^{\prime},\theta_{2}^{\prime}).$
Then,
$\mathbf{x}_{1}^{*}$ $\sim\theta_{1}^{\prime}Dir(\alpha_{11}^{\prime},\alpha_{12}^{\prime},\alpha_{11}^{\prime},\alpha_{12}^{\prime})+\theta_{2}^{\prime}Dir(\alpha_{21}^{\prime},\alpha_{22}^{\prime},\alpha_{21}^{\prime},\alpha_{22}^{\prime})$ (17)
$=\theta_{2}Dir(\alpha_{21},\alpha_{22},\alpha_{21},\alpha_{22})+\theta_{1}Dir(\alpha_{11},\alpha_{12},\alpha_{11},\alpha_{12}),$
which will clearly give rise to the same likelihood value as the originally ordered parameters.
To address this, we create a hard partition based on probabilities $\hat{\mathbf{Z}}$ from the E-step and use the order of clusters in this hard partition as reference since column entries of $\hat{\mathbf{Z}}$ are tied to the individual terms of the Dirichlet mixture. Then we use the $matchClasses()$ function in R ($e1071$ package, Dimitriadou et al.,, 2009) to match the order of clusters in the E-step to that in the C-step, and use this match to reorder the rows of $\hat{\mathbf{A}}$ and entries of $\hat{\boldsymbol{\theta}}$ at the convergence of the algorithm.
@@ 3.3 Algorithm initialisation
As in most mixture models, extra attention should be paid to the choice of the initialisation strategy since the objective function in question is not unimodal, and the algorithm is only guaranteed to converge to a local optimum (Wu,, 1983).
In this paper, we consider the following initialisation strategies, which are compared in the simulation study in Section 5.1:
@@ 3.4 Model selection
One popular choice in stochastic block modelling is the integrated completed likelihood criterion (ICL, Biernacki et al.,, 2000; Daudin et al.,, 2008), typically employed for selecting the number of groups in this context; see for example Rastelli and Fop, (2020), Ng and Murphy, (2021), and El Haj et al., (2022). In the case of DirSBM, the ICL is given by
@@
where $l^{hyb}_{c}(\mathbf{\hat{A}},\boldsymbol{\hat{\theta}}|\mathbf{X},\mathbf{\hat{Z}},K)$ is the complete data hybrid log-likelihood from Equation (11), evaluated using parameter estimates at convergence of the algorithm, used in place of the complete data log-likelihood following Marino and Pandolfi, (2022). We fit the DirSBM with different numbers of clusters, and compute the respective values of the criterion. The highest ICL value corresponds to the optimal number of groups.
Appendix C contains derivations of the ICL that follow closely the derivations of the original ICL for stochastic block models proposed in Daudin et al., (2008). The simulation study concerning model selection performance using the proposed framework is contained in Section 5.4.
@@ 4 An alternative approach
The centered log-ratio transformation has been chosen for its property of preserving the dimensionality of the observations. That is, it maps the Aitchison simplex sample space $\mathbb{S}^{n}$ to the standard Euclidean space $\mathbb{R}^{n}$, unlike many other transformations that map to $\mathbb{R}^{n-1}$ (Greenacre,, 2021). In this case, each dimension of $\mathbf{u}_{i}$ corresponds to a transformed weight of an edge connecting node $i$ to each of the other nodes in the network, so the reduction of dimensionality would require a choice of an edge to be dropped, leading to information loss.
@@
There are two major drawbacks of the alternative approach in comparison to the DirSBM. First of all, the normality assumption in Equation (20) may not be appropriate as the mapping of compositional vectors to the real line (whether by using CLR or any other transformation) only eliminates a unit-sum constraint and does not guarantee convenient distributional behaviour of the resulting transformed data. This means that, in order to utilise this approach, one needs to check the appropriateness of the transformation and the type of weighted SBM to fit to the transformed data. The DirSBM, on the other hand, is a ready-to-use model that takes into account the unique behaviour of the compositional data as it models them directly. The second disadvantage of the transformation-based approach is interpretability. The parameters of the distribution in Equation (20) (or of any other distribution selected to model the transformed data) refer to the auxiliary Euclidean space, and it is unclear how one could use these to understand the patterns present in the original simplex space. As presented in Section 2.1.1, the parameters of the DirSBM can be used to compute the expected exchange proportions between nodes as well as clusters, making the interpretation of results intuitive.
@@ 5 Simulation studies
In this section, we consider various simulation studies to assess the performance of our model with respect to clustering, parameter estimation, and model selection. We also investigate the effect of the different initialisation strategies listed in Section 3.3 on the quality of the final inferred clustering of the nodes.
Artificial data are generated according to the DirSBM presented in Section 2. Different numbers of clusters, $K=\{2,3,5\}$, and network sizes, $n=\{30,50,70,100\}$ are considered. The parameter matrices $\mathbf{A}$ are defined so that the entries have varying levels of homogeneity, resulting in indirect modifications to the degree of overlap between clusters in the compositions. The term “level of homogeneity” is used to describe the extent to which the between-cluster weight parameters, i.e. $\{\alpha_{kh}\}_{h\neq k}$, are close to the respective within-cluster weight parameters, $\alpha_{kk}$. When the values of the between- and the within-cluster parameters are set to be more similar, the resulting compositional weights in the network become more homogeneous, thus making separating the clusters more challenging. On the other hand, the lower the homogeneity, the more separated the clusters. Further details on the parameter values used in the simulation studies can be found in Appendix D.
Throughout this section, where relevant, clustering performance is evaluated using the adjusted Rand Index (ARI, Rand,, 1971; Hubert and Arabie,, 1985), which measures the agreement between two partitions, adjusted for random chance, by comparing the estimated clustering against the true classification of the nodes.
@@ 5.1 Choice of initialisation strategy
Figure 1: Adjusted Rand Index (ARI) of DirSBM with different initialisation strategies with mean ARI in blue.
We initialise the algorithm using the strategies outlined in Section 3.3 and compute the ARI. The results of this simulation study are shown in Figure 1. From the figure, random initialisation seems to work best on average, with a mean ARI of $0.805$, despite being the only strategy that is not informed by the data. It also exhibits relatively low variability in comparison to all but one other strategy, which is k-means on log-ratio transformed data that is clearly inferior, with a mean ARI of only $0.513$. Further investigations also reveal that, as well as having the best average $ARI$ and low variability, random initialisation has led to the highest observed hybrid log-likelihood value in 33 out of 50 instances. Given that the number of random partitions considered was only 5, this result also shows that the proposed algorithm is capable of recovering the clustering patterns fairly successfully even when the initial cluster labels are assigned randomly and are not based on some educated guess.
It may come as a surprise to the reader that out of all the strategies considered, random initialisation performed best, as it gives a starting cluster partition that is completely uninformed. However, there are two aspects of the data modelled by the DirSBM that are worth remembering. Firstly, the data matrix $\mathbf{X}^{*}$ introduced in Section 2 is quite unusual in its structure in the sense that the ordering of the dimensions in the compositional edge weight vectors is not fixed and is tied to the index of the sender node. For example, the first entry of observation $\mathbf{x}^{*}_{1}$, i.e. $x^{*}_{11}$, refers to the weight of an edge from node 1 to node 2 (as there are no self-loops and so there is no weight corresponding to an edge from node $1$ to itself), whereas, for any other node $i\neq 1$, $x^{*}_{i1}$ is the edge weight from node $i$ to node 1. The same shift along the dimensions happens across the whole set of compositional weight vectors as we take each node in turn to be the sender. This means that distance-based algorithms, such as k-means, would not necessarily compute the distances between the matching pairs of nodes’ edge weights, e.g. the first set of distances for node $1$ would involve the edge weight from node 1 to node 2, and for any other node $i$, it would involve the edge weight from node $i$ to node 1. This can potentially lead to finding spurious clustering patterns, producing initial partitions that result in convergence to local maxima.
@@
Throughout the simulation studies, we use random initialisation with 5 starts, and to increase the reliability of results for the real world data in sections 6.1 and 6.2 we increase the number of random partitions to 20.
@@ 5.2 Parameter estimation performance
Figure 2: Frobenius distance between the true parameter matrix $\mathbf{A}$ and its estimate $\mathbf{\hat{A}}$, based on 50 artificial data sets. The number of clusters $K$ is fixed at the respective true values of 2(a) 2, 2(b) 3 and 2(c) 5.
@@ 5.3 Clustering structure recovery
Figure 3: Adjusted Rand Index between the true partition and clustering solutions estimated by binary SBM, Gaussian SBM, Gaussian SBM on centered log-ratio transformed data and DirSBM, with different network sizes and levels of Dirichlet parameters homogeneity. Results are based on 50 artificial data sets with 3(a) $K=2$, 3(b) $K=3$ and 3(c) $K=5$ clusters.
@@
Figure 3 illustrates the performance of all 4 models on artificial data sets with 2, 3 and 5 clusters across multiple network sizes and varying levels of Dirichlet parameter homogeneity. Each boxplot is based on 50 networks. The DirSBM was initialised with 5 random cluster partitions based on the results of the simulation study in Section 5.1.
It is evident that the performance of all models improves with sample size, often approaching an ARI of $1$ as we reach 100 nodes. In the case of BinSBM and GausSBM, this can be due to the fact that, in larger networks, the compositional samples are rather high-dimensional, and so the constant-sum constraint does not induce as much dependence between the individual dimensions as in smaller networks, resulting in a good fit of the models that assume independence between the weights. In the case of CLR+GausSBM and DirSBM, the improvement of performance with sample size can be associated with more accurate parameter estimation. Smaller networks present a greater challenge, due to both stronger dependence between the dimensions of compositions and the limited amount of data available for parameter estimation.
The CLR+GausSBM and DirSBM tend to outperform the competitors in terms of clustering structure recovery, and the gap in performance is significantly wider for smaller networks and higher homogeneity of Dirichlet parameters. It is also notable that the performance of BinSBM and the GausSBM on the raw compositional data often exhibits much higher variability indicated by wider boxplots.
With regards to the comparison between the DirSBM and the CLR+GausSBM, the two methods tend to perform similarly well. Looking at Figures 3(b) and 3(c), we observe that in many scenarios the DirSBM performs better and with lower overall variability, making it a more a suitable model in the cases with larger number of clusters in the data. Considering Figure 3(a), where the number of underlying clusters is 2, we note that CLR+GausSBM tends to perform better than DirSBM. Albeit a simplistic approach, CLR+GausSBM seems to present a valid alternative when modelling compositional networks if it is expected that the number of clusters is small.

### v2

@@ 3.2 Identifiability of parameter ordering
The DirSBM exhibits a minor non-idenfitiability issue related to the ordering of clusters and the ordering of parameters. In particular, once we obtain the final node partition and the parameter estimates using the algorithm described in Section 3.1, the ordering of rows of the estimated Dirichlet concentration matrix $\hat{\mathbf{A}}$ and of the entries of the mixing proportions vector $\hat{\boldsymbol{\theta}}$ is not guaranteed to match the ordering of clusters in the partition. To aid understanding, consider an example with
$\mathbf{c}=(1,1,2,1,2),\leavevmode\nobreak\ \mathbf{A}=\begin{pmatrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{pmatrix},\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \boldsymbol{\theta}=(\theta_{1},\theta_{2}).$
For the first observation, which is placed in cluster 1, $\mathbf{x}_{1}^{*}\sim\theta_{1}Dir(\alpha_{11},\alpha_{12},\alpha_{11},\alpha_{12})+\theta_{2}Dir(\alpha_{21},\alpha_{22},\alpha_{21},\alpha_{22})$. Now, consider the same partition of nodes with the same cluster order and permute the rows of the parameter matrix $\mathbf{A}$ and the entries of $\boldsymbol{\theta}$, i.e.
$\mathbf{c}=(1,1,2,1,2),\leavevmode\nobreak\ \mathbf{A}^{\prime}=\begin{pmatrix}\alpha_{21}&\alpha_{22}\\ \alpha_{11}&\alpha_{12}\end{pmatrix}=\begin{pmatrix}\alpha_{11}^{\prime}&\alpha_{12}^{\prime}\\ \alpha_{21}^{\prime}&\alpha_{22}^{\prime}\end{pmatrix},\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \boldsymbol{\theta}^{\prime}=(\theta_{2},\theta_{1})=(\theta_{1}^{\prime},\theta_{2}^{\prime}).$
Under the permutation of the parameters $\mathbf{A}$ and $\boldsymbol{\theta}$, the first observation is distributed $\mathbf{x}_{1}^{*}\sim\theta_{1}^{\prime}Dir(\alpha_{11}^{\prime},\alpha_{12}^{\prime},\alpha_{11}^{\prime},\alpha_{12}^{\prime})+\theta_{2}^{\prime}Dir(\alpha_{21}^{\prime},\alpha_{22}^{\prime},\alpha_{21}^{\prime},\alpha_{22}^{\prime})$. Clearly:
$\theta_{1}^{\prime}D(\mathbf{x}_{1}^{*};\alpha_{11}^{\prime},\alpha_{12}^{\prime},\alpha_{11}^{\prime},\alpha_{12}^{\prime})+\theta_{2}^{\prime}D(\mathbf{x}_{1}^{*};\alpha_{21}^{\prime},\alpha_{22}^{\prime},\alpha_{21}^{\prime},\alpha_{22}^{\prime})$
$=\theta_{2}D(\mathbf{x}_{1}^{*};\alpha_{21},\alpha_{22},\alpha_{21},\alpha_{22})+\theta_{1}D(\mathbf{x}_{1}^{*};\alpha_{11},\alpha_{12},\alpha_{11},\alpha_{12})$
where, for ease of notation, $D(\mathbf{x};\ldots)$ denotes the Dirichlet density evaluated at $\mathbf{x}$ with the corresponding concentration parameters. Therefore, the permuted parameters $\mathbf{A}^{\prime}$ and $\boldsymbol{\theta}^{\prime}$ yield the same likelihood value as the parameters in the original order.
To address this issue, we create a hard partition based on the estimated probabilities $\hat{\mathbf{Z}}$ from the E-step. Subsequently, we use the order of clusters in this hard partition as reference since the column entries of $\hat{\mathbf{Z}}$ are tied to the individual terms of the Dirichlet mixture. Lastly, we use the $matchClasses()$ function in R ($e1071$ package, Dimitriadou et al.,, 2009) to match the order of clusters in the E-step to that in the C-step, and use this match to reorder the rows of $\hat{\mathbf{A}}$ and entries of $\hat{\boldsymbol{\theta}}$ at the convergence of the algorithm.
@@ 3.3 Algorithm initialisation
As in most mixture models, extra attention should be paid to the choice of the initialisation strategy since the objective function in question is not unimodal, and the algorithm is only guaranteed to converge to a local optimum (Wu,, 1983). In this paper, we consider the following initialisation strategies, which are compared in the simulation study in Section 5.1:
@@ 3.4 Model selection
A popular choice in stochastic block modelling is the integrated completed likelihood criterion (ICL, Biernacki et al.,, 2000; Daudin et al.,, 2008), typically employed for selecting the number of groups in this context; see for example Rastelli and Fop, (2020), Ng and Murphy, (2021), and El Haj et al., (2022). In the case of DirSBM, the ICL is given by
@@
where $l^{hyb}_{c}(\mathbf{\hat{A}},\boldsymbol{\hat{\theta}}|\mathbf{X},\mathbf{\hat{Z}},K)$ is the complete data hybrid log-likelihood from Equation (11), evaluated using parameter estimates at convergence of the algorithm, used in place of the complete data log-likelihood following Marino and Pandolfi, (2022). We fit the DirSBM with different numbers of clusters, and compute the respective values of the criterion. The highest ICL value corresponds to the optimal number of groups. C contains derivations of the ICL that follow closely the derivations of the original ICL for stochastic block models proposed in Daudin et al., (2008). The simulation study concerning model selection performance using the proposed criterion is contained in Section 5.4.
@@ 4 An alternative approach
The centered log-ratio transformation has been chosen for its property of preserving the dimensionality of the observations. That is, it maps the Aitchison simplex sample space $\mathbb{S}^{n}$ to the standard Euclidean space $\mathbb{R}^{n}$, unlike many other transformations that map to $\mathbb{R}^{n-1}$ (Greenacre,, 2021). In this case, each dimension of $\mathbf{u}_{i}$ corresponds to a transformed weight of an edge connecting node $i$ to each of the other nodes in the network, hence the reduction of dimensionality would require a choice of an edge to be dropped, leading to information loss.
@@
There are two major drawbacks of the alternative approach in comparison to the DirSBM. First of all, the normality assumption in Equation (19) may not be appropriate as the mapping of compositional vectors to the real line (whether by using CLR or any other transformation) only eliminates a unit-sum constraint and does not guarantee convenient distributional behaviour of the resulting transformed data. This means that, in order to utilise this approach, one needs to check the appropriateness of the transformation and the type of weighted SBM to fit to the transformed data. The DirSBM, on the other hand, is a ready-to-use model that takes into account the unique behaviour of the compositional data as it models them directly. The second disadvantage of the transformation-based approach is interpretability. The parameters of the distribution in Equation (19) (or of any other distribution selected to model the transformed data) refer to the auxiliary Euclidean space, and it is unclear how one could use these to understand the patterns present in the original simplex space. As presented in Section 2.2, the parameters of the DirSBM can be used to compute the expected exchange proportions between nodes as well as clusters, making the interpretation of results intuitive.
@@ 5 Simulation studies
In this section, we consider various simulation studies to assess the performance of our model with respect to clustering, parameter estimation, and model selection. We also investigate the effect of the different initialisation strategies listed in Section 3.3 on the quality of the final inferred clustering of the nodes. Additionally, we examine the effect on the model’s performance of the presence of zero-weighted edges in the network. Throughout this section, where relevant, clustering performance is evaluated using the adjusted Rand index (ARI, Rand,, 1971; Hubert and Arabie,, 1985), which measures the agreement between two partitions, adjusted for random chance, by comparing the estimated clustering against the true classification of the nodes.
Artificial data are generated according to the DirSBM presented in Section 2. Different numbers of clusters, $K=\{2,3,5\}$, and network sizes, $n=\{30,50,70,100\}$ are considered. The parameter matrices $\mathbf{A}$ are defined so that the entries have varying levels of homogeneity, resulting in indirect modifications to the degree of overlap between clusters in the compositions. The term “level of homogeneity” is used to describe the extent to which the between-cluster weight parameters, i.e. $\{\alpha_{kh}\}_{h\neq k}$, are close to the respective within-cluster weight parameters, $\alpha_{kk}$. When the values of the between- and the within-cluster parameters are set to be more similar, the resulting compositional weights in the network become more homogeneous, thus making separating the clusters more challenging. On the other hand, the lower the homogeneity, the more separated the clusters. Further details on the parameter values used in the simulation studies can be found in D.
@@ 5.1 Choice of initialisation strategy
Figure 1: Adjusted Rand index (ARI) of DirSBM with different initialisation strategies, with the mean value of ARI indicated in blue.
We initialise the algorithm using the strategies outlined in Section 3.3 and compute the ARI. The results of this simulation study are shown in Figure 1. From the figure, random initialisation seems to work best on average, with a mean ARI of $0.805$, despite being the only strategy that is not informed by the data. It also exhibits relatively low variability in comparison to all but one other strategy, which is k-means on log-ratio transformed data that is clearly inferior, with a mean ARI of only $0.513$. Further investigations also reveal that, as well as having the best average $ARI$ and low variability, random initialisation has led to the highest observed hybrid log-likelihood value in 33 out of 50 instances. Given that the number of random partitions considered was only 5, this result also shows that the proposed algorithm is capable of recovering the clustering patterns fairly successfully even when the initial cluster labels are assigned randomly and are not based on some educated guess. Throughout the simulation studies, we use random initialisation with 5 starts, and to increase the reliability of results for the real world data in Section 6 we increase the number of initial random partitions to 20.
It may come as a surprise to the reader that out of all the strategies considered, random initialisation performed best, as it gives a starting cluster partition that is completely uninformed. However, there are two aspects of the data modelled by the DirSBM that are worth remembering. Firstly, the data matrix $\mathbf{X}^{*}$ introduced in Section 2 is quite unusual in its structure in the sense that the ordering of the dimensions in the compositional edge weight vectors is not fixed and is tied to the index of the sender node. For example, the first entry of observation $\mathbf{x}^{*}_{1}$, i.e. $x^{*}_{11}$, refers to the weight of an edge from node 1 to node 2 (as there are no self-loops and hence there is no weight corresponding to an edge from node $1$ to itself), whereas, for any other node $i\neq 1$, $x^{*}_{i1}$ is the edge weight from node $i$ to node 1. The same shift along the dimensions happens across the whole set of compositional weight vectors as we take each node in turn to be the sender. This means that distance-based algorithms, such as k-means, would not necessarily compute the distances between the matching pairs of nodes’ edge weights, e.g. the first set of distances for node $1$ would involve the edge weight from node 1 to node 2, and for any other node $i$, it would involve the edge weight from node $i$ to node 1. This can potentially lead to finding spurious clustering patterns, producing initial partitions that result in convergence to local maxima.
@@ 5.2 Parameter estimation performance
Figure 2: Frobenius distance between the true parameter matrix $\mathbf{A}$ and its estimate $\mathbf{\hat{A}}$, based on 50 artificial data sets for varying levels of Dirichlet parameters homogeneity. The number of clusters $K$ is fixed at the respective true values of 2(a) 2, 2(b) 3 and 2(c) 5.
@@ 5.3 Clustering structure recovery
Figure 3: Adjusted Rand index (ARI) between the true partition and clustering solutions estimated by binary SBM, Gaussian SBM, Gaussian SBM on centered log-ratio transformed data and DirSBM, with different network sizes $(n)$ and levels of Dirichlet parameters homogeneity. Results are based on 50 artificial data sets with 3(a) $K=2$, 3(b) $K=3$ and 3(c) $K=5$ clusters.
@@
Figure 3 illustrates the performance of all 4 models on artificial data sets with 2, 3 and 5 clusters across multiple network sizes and varying levels of Dirichlet parameter homogeneity. Each boxplot is based on 50 networks. The DirSBM was initialised with 5 random cluster partitions based on the results of the simulation study in Section 5.1. It is evident that the performance of all models improves with sample size, often approaching an ARI of $1$ as we reach 100 nodes. In the case of BinSBM and GausSBM, this can be due to the fact that, in larger networks, the compositional samples are rather high-dimensional, hence the constant-sum constraint does not induce as much dependence between the individual dimensions as in smaller networks, resulting in a good fit of the models that assume independence between the weights. In the case of CLR+GausSBM and DirSBM, the improvement of performance with sample size can be associated with more accurate parameter estimation. Smaller networks present a greater challenge, due to both stronger dependence between the dimensions of compositions and the limited amount of data available for parameter estimation.
The CLR+GausSBM and DirSBM tend to outperform the competitors in terms of clustering structure recovery, and the gap in performance is significantly wider for smaller networks and higher homogeneity of Dirichlet parameters. It is also notable that the performance of BinSBM and the GausSBM on the raw compositional data often exhibits much higher variability indicated by wider boxplots. With regards to the comparison between the DirSBM and the CLR+GausSBM, the two methods tend to perform similarly well. Looking at Figures 3(b) and 3(c), we observe that in many scenarios the DirSBM performs better and with lower overall variability, making it a more a suitable model in the cases with larger number of clusters in the data. Considering Figure 3(a), where the number of underlying clusters is 2, we note that CLR+GausSBM tends to perform better than DirSBM. Albeit a simplistic approach, CLR+GausSBM seems to present a valid alternative when modelling compositional networks if it is expected that the number of clusters is small.